# Jet schemes, Newton polygons and continued fractions

### v2

@@ 5. Jet Schemes of Newton Plane Curves
Let $f=f_{1}...f_{r}\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate product of $r$ branches and let $g=g_{1}...g_{r}$ be another such polynomial with the same number of branches. Assume
@@ 6. Useful Results
In this part of the article, we will mention the key results that will be useful in proving the theorems in the next sections.
For $\alpha\in\mathbf{R}^{2}$, set $|\alpha|=\alpha_{1}+\alpha_{2}$.
Moreover, we say that two polynomials $f,g\in k[x_{1},...,x_{n}]$ are congruent modulo a variety $W\subset\mathbf{A}^{\infty}$ (which we write $f=g\mod W$) if they are congruent modulo the defining ideal of $W$ in $k[x_{1},x_{2},$...].
This allows us to introduce the terminology, $f=g\mod\alpha$ when $W=\mathcal{H}^{\alpha}_{\infty}$. Let’s start with this observation which will be used later.
@@
Write $in_{\alpha}(f)=y^{p}-cx^{q}$. Then,
$\nu_{\alpha}(f)=spq>s(p+q)=|\alpha|$
@@
The below lemma will be used (in the next section) in proving Theorem 7.
@@
$F^{\alpha}_{m}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}$
@@
    and since $T(f)$ is the semi-line that passes by $(0,0)$ and $\alpha$ then for all $i$:
@@
    which isn’t a monomial. Moreover, $in_{\alpha}(f_{j})$ are pairwise coprime as polynomials since $f$ is Newton non-degenerate. This allows to say that:
    $F^{\alpha}_{\nu_{\alpha}(f)}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})}$
    This proves the initial case $m=\nu_{\alpha}(f)$.
@@
    $F^{\alpha}_{m-1}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-1-\nu_{\alpha}(f)}$
    Then, as $F^{\alpha}_{m}=F^{\alpha}_{m-1}\cap\{f^{(m)}=0\}$, let’s show that:
    $F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-1-\nu_{\alpha}(f)}\cap\{f^{(m)}=0\}=F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}$
    This comes from the fact that if we set $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$ then
    $f^{(m)}=f^{(k)}_{i}\prod_{j\neq i}f^{[\nu_{\alpha}(f_{i})]}_{i}\mod F^{\alpha}_{i,k-1}$
    So the proposition is true for $m$.
This concludes the induction. ∎
We will show (in the next section) that the irreducible components of the jet schemes $\mathcal{C}^{0}_{m}$ (over the origin) of Newton plane curves $\mathcal{C}=\{f=0\}\subset\mathbf{A}^{2}$ are of two kinds:
  • the ones of the form $\mathcal{H}^{\alpha}_{m}$ who are irreducible and have $|\alpha|$ as codimension. We take the $\alpha$ such that there is no $\beta\leq\alpha$ component-wise with $\alpha\neq\beta$. These $\mathcal{H}^{\alpha}_{m}$ are called having hyperplane coordinates;
@@
The next lemma will allow us to have their codimensions and verify their irreducibility.
@@
Denote by $\pi:\mathbf{A}^{2}_{m}\rightarrow\mathbf{A}^{2}_{k}$ the truncation morphism of the jet schemes of $\mathbf{A}^{2}$ : they are defined as $\pi(x_{0},...,x_{m},y_{0},...,y_{m})=(x_{0},...,x_{k},y_{0},...,y_{k})$ where $v_{i}=v^{(i)}$ if $v=x$ or $v=y$. Notice that:
$F^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}=\pi^{-1}(F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k})$
So to have $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}$ irreducible, it’s enough to have $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k}$ irreducible. Since $f$ defines a Newton plane branch, then $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k}$ is irreducible by Lemma 4.3 of [14] and using our notations it has codimension $|\alpha|+k-\nu_{\alpha}(f)+1$. If we lift to level $m$ the irreducibility is preserved as well as the codimension. ∎
Now that we know who are the candidates to be irreducible components of $\mathcal{C}^{0}_{m}$ for all $m$, we need to compare (by inclusion) these candidates to be certain that they have the property of being components. Their irreducibility has been proven in the previous result.
First, let’s recall some notations: let $\mathbf{A}^{2}_{m}$ be the $m$-jet space of $\mathbf{A}^{2}$ with variables $x^{(0)},...,x^{(m)},y^{(0)},...,y^{(m)}$. Let $\alpha\in\mathbf{N}^{2}$.
  • Set $\mathcal{H}^{\alpha}_{m}=\{x^{(i)}=y^{(j)}=0:0\leq i<\alpha_{1},0\leq j<\alpha_{2}\}\subset\mathbf{A}^{2}_{m}$;
  • Let $g$ define a branch in $\mathbf{A}^{2}$ and $\nu_{\alpha}(g)\leq k\leq m$, set
    $\mathbf{A}^{2}_{m}\cap G^{\alpha}_{k}=\{x^{(\alpha_{1})}\neq 0\}\cap\{y^{(\alpha_{2})}\neq 0\}\cap\mathcal{H}^{\alpha}_{m}\cap\{g^{[\nu_{\alpha}(g)]}=...=g^{(k)}=0\}$
Let’s sum up the purpose of this section by Table 2: whenever $\mathcal{H}^{\gamma}_{m}$ is mentioned we impose the condition $\nu_{\gamma}(f)>m$ and whenever
@@
corresponds to a branch defined by a factor $f_{i}$ of $f$ with $\alpha\in T(f_{i})$, we impose $m\geq\nu_{\alpha}(f)$ and $a=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$.
AA $B$ when does $A\subseteq B?$ why?
@@
$\mathcal{H}^{\gamma}_{m}$ $\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,a}}$ $\gamma=\alpha+(1,1)$, $m=\nu_{\alpha}(f)$
@@
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,a}}$ $\overline{\mathbf{A}^{2}_{m}\cap F^{\beta}_{i,b}}$ $\alpha=\beta$ and $f_{i}=f_{j}$ compare codimensions
@@
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve and $f_{i}$ define branch of $f$. Write $T(f_{i})=\mathbf{R}_{\geq 0}(p,q)$. Let $\alpha,\beta\in T(f_{i})\cap\mathbf{N}^{2}$ and set $\gamma=\beta+(1,1)$ and suppose $\nu_{\gamma}(f)>m\geq\nu_{\alpha}(f)$. Then, we have
@@
if and only if $\alpha=\beta$ and $m=\nu_{\alpha}(f)$. Moreovoer, this inclusion is always strict.
@@
Suppose first $\alpha=\beta$ and that: $\mathcal{H}^{\gamma}_{m}=\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$. So,
$U:=F^{\alpha}_{i,\nu_{\alpha}(f_{i})+1}\cap\mathbf{A}^{2}_{m}\subset\mathcal{H}^{\gamma}_{m}$
However, $x^{(\alpha_{1})}\neq 0$ in $U$ and $x^{(\alpha_{1})}=0$ in $\mathcal{H}^{\gamma}_{m}$, which is a contradiction. So the inclusion if it occurs is always strict. Now, let’s show the equivalence in the theorem: Suppose that $m=\nu_{\alpha}(f)$ and $\alpha=\beta$. Then,
@@
with $f^{[\nu_{\alpha}(f_{i})]}_{i}=[{y^{(\alpha_{2})}}]^{p}-c_{i}\cdot[{x^{(\alpha_{1})}}]^{q}\mod\alpha$ where $T(f_{i})=\mathbf{R}_{\geq 0}(p,q)$. Then, we have the inclusion: $\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$, with $m=\nu_{\alpha}(f)$ and $\gamma=\alpha+(1,1)$.
@@
Now given that we have the inclusion above with $m\geq\nu_{\alpha}(f)$ together with the condition $\alpha=\beta$. We have:
$\mathcal{H}^{\gamma}_{m}\neq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$
Let’s compare the codimensions of these irreducible sets:
$|\gamma|=|\alpha|+2>|\alpha|+1+m-\nu_{\alpha}(f)$
Then, $1>m-\nu_{\alpha}(f)\geq 0$. This shows $m=\nu_{\alpha}(f)$.
@@
Now, assume on the other hand $\alpha\neq\beta$ with $\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$. So, $\mathcal{H}^{\gamma}_{m}\subseteq\mathcal{H}^{\alpha}_{m}$ and $\gamma\geq\alpha$ component-wise. By comparing codimensions,
$|\gamma|\geq 1+|\alpha|+m-\nu_{\alpha}(f)$
Write $\beta=b(p,q)$ and $\alpha=a(p,q)$ then $\gamma=(bp+1,bq+1)$. Using $\gamma\geq\alpha$ we find $bp+1\geq ap$ and $bq+1\geq aq$. Since $\alpha\neq\beta$ then $a\neq b$. If $a>b$ then
$bq+1\geq aq>bq$
are integers. Then $aq=bq+1$ then (ab)q=1(a-b)q=1. So, q=ab=1q=a-b=1. Using $bp+1\geq ap>bp$ we get p=q=ab=1p=q=a-b=1. So, $\beta=(b,b)$ and $\alpha=(b+1,b+1)$. Thus, $\gamma=\beta+(1,1)=\alpha$. Now, recall $\nu_{\gamma}(f)>m\geq\nu_{\alpha}(f)$ this a contradiction with $\gamma=\alpha$. So, if $\alpha\neq\beta$ then
$\mathcal{H}^{\gamma}_{m}\not\subset\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$
Now suppose $\alpha\neq\beta$ with $m=\nu_{\alpha}(f)$. If we had
$\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})}\cap\mathbf{A}^{2}_{m}}\subseteq\mathcal{H}^{\alpha}_{m}$
Then by comparing codimension, $|\gamma|\geq 1+|\alpha|$. Also we have component-wise $\gamma\geq\alpha$. Write again $\gamma=(bp+1,bq+1)$ and $\alpha=(ap,aq)$. Thus,
$b(p+q)+2\geq 1+a(p+q)$
and $b(p+q)+1\geq a(p+q)$. Since $a\neq b$ and if $a>b$ then
$b(p+q)+1\geq a(p+q)>b(p+q)$
Since we are working with integers then (ab)(p+q)=1(a-b)(p+q)=1. Thus,
ab=p=q=1a-b=p=q=1
As before, we get $\gamma=\alpha$ and a contradiction with $\nu_{\gamma}(f)>m\geq\nu_{\alpha}(f)$. ∎
@@
Now, let’s make sure (in the next theorem) that distinct infinite components (of the jet schemes of a Newton Plane Curves) come from distinct branches of this curve.
In other terms, if one such irreducible component is inside another infinite component then these components are in fact the same.
@@
Before that, recall the formula of the codimension of $G^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}$ where $G$ is defined by an irreducible Newton non-degenerate $g$ that divides $f$
@@
where we set $k=\nu_{\alpha}(g)+m-\nu_{\alpha}(f)$.
@@
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve. Let $f_{i}$ and $f_{j}$ define branches of $f$ such that
$T(f_{l})=\mathbf{R}_{\geq 0}\alpha_{l}:\alpha_{l}\in\mathbf{N}^{2}$
and $\nu_{\alpha_{l}}(f)\leq m$ for both $l\in\{i,j\}$. Set $k_{l}=\nu_{\alpha_{l}}(f_{l})+m-\nu_{\alpha_{l}}(f)$. If
@@
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{j}}_{j,k_{j}}}$

### v3

@@ 5. Jet Schemes of Newton Plane Curves
Let $f=f_{1}...f_{r}\in\mathcal{O}_{k^{2}}$ be a Newton non-degenerate product of $r$ branches and let $g=g_{1}...g_{r}$ be another such polynomial with the same number of branches. Assume
@@ 6. Useful Results
In this part of the article, we shall present the key results that are very useful in proving our theorems in the remaining sections.
We say that two polynomials $f,g\in k[x_{1},...,x_{n}]$ are congruent modulo a variety $W\subset\mathbf{A}^{\infty}$ (which we write $f=g\mod W$) if they are congruent modulo the defining ideal of $W$ in $k[x_{1},x_{2},$...].
Now for $\alpha\in\mathbf{R}^{2}$, we set $|\alpha|=\alpha_{1}+\alpha_{2}$. With this in mind, we can introduce the terminology, $f=g\mod\alpha$ when $W=\mathcal{H}^{\alpha}_{\infty}$. But first we need the following observation which will be useful for later.
@@
Write $in_{\alpha}(f)=y^{p}-cx^{q}$. Then, $pq>p+q$ since $p,q>1$.
$\nu_{\alpha}(f)=spq>s(p+q)=|\alpha|.$
@@
The next lemma will be used (in the next section) in proving Theorem 7.
@@
$F^{\alpha}_{m}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}.$
@@
    and since $T(f)$ is the semi-line that passes through $(0,0)$ and $\alpha$, then for all $i$ it holds that
@@
    which is clearly not a monomial. Moreover, the $in_{\alpha}(f_{j})$ are pairwise coprime as polynomials since $f$ is Newton non-degenerate. This allows us to conclude that:
    $F^{\alpha}_{\nu_{\alpha}(f)}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})},$
    which proves the initial case $m=\nu_{\alpha}(f)$.
@@
    $F^{\alpha}_{m-1}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-1-\nu_{\alpha}(f)}.$
    Then, as $F^{\alpha}_{m}=F^{\alpha}_{m-1}\cap\{f^{(m)}=0\}$, we need to show that:
    $F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-1-\nu_{\alpha}(f)}\cap\{f^{(m)}=0\}=F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}.$
    However, this can be seen from the fact that if $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$ then
    $f^{(m)}=f^{(k)}_{i}\prod_{j\neq i}f^{[\nu_{\alpha}(f_{i})]}_{i}\mod F^{\alpha}_{i,k-1}.$
    So the proposition is true for $m$, and this concludes the induction.
In the next section, we shall show that the irreducible components of the jet schemes $\mathcal{C}^{0}_{m}$ (over the origin) of Newton plane curve singularities $\mathcal{C}=\{f=0\}\subset\mathbf{A}^{2}$ are of two kinds:
  • the ones of the form $\mathcal{H}^{\alpha}_{m}$ that are irreducible and having $|\alpha|$ as codimension. In this case, the $\alpha$’s are chosen so that there is no $\beta\leq\alpha$ component-wise with $\alpha\neq\beta$. These $\mathcal{H}^{\alpha}_{m}$ are said to have hyperplane coordinates;
@@
The next lemma is concerned with determining their codimensions and verifying their irreducibility.
@@
Denote by $\pi:\mathbf{A}^{2}_{m}\rightarrow\mathbf{A}^{2}_{k}$ the truncation morphism of the jet schemes of $\mathbf{A}^{2}$. They are defined as $\pi(x_{0},...,x_{m},y_{0},...,y_{m})=(x_{0},...,x_{k},y_{0},...,y_{k})$ where $v_{i}=v^{(i)}$ if $v=x$ or $v=y$. Notice that
$F^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}=\pi^{-1}(F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k}).$
In order to have $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}$ irreducible, it’s enough to have $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k}$ irreducible. Since $f$ defines a Newton plane branch, then $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k}$ is irreducible by Lemma 4.3 of [14] and using our notation it has codimension $|\alpha|+k-\nu_{\alpha}(f)+1$. Note that at the level $m$, the irreducibility of $F_{k}^{\alpha}\cap\mathbf{A}^{2}_{m}$ is preserved as well as the codimension since it has the same defining equations as $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k}$. ∎
To distinguish the candidates for the irreducible components of $\mathcal{C}^{0}_{m}$ for all $m$, we only need to compare them (by inclusion) as their irreducibility has been proven in the previous result.
First, let us remind the reader with our notation: $\mathbf{A}^{2}_{m}$ is the $m$-jet space of $\mathbf{A}^{2}$ with variables $x^{(0)},...,x^{(m)},y^{(0)},...,y^{(m)}$ and for $\alpha\in\mathbf{N}^{2}$,
  • $\mathcal{H}^{\alpha}_{m}=\{x^{(i)}=y^{(j)}=0:0\leq i<\alpha_{1},0\leq j<\alpha_{2}\}\subset\mathbf{A}^{2}_{m}$;
  • Let $g$ define a branch in $\mathbf{A}^{2}$ and $\nu_{\alpha}(g)\leq k\leq m$, Next, set
    $\mathbf{A}^{2}_{m}\cap G^{\alpha}_{k}=\{x^{(\alpha_{1})}\neq 0\}\cap\{y^{(\alpha_{2})}\neq 0\}\cap\mathcal{H}^{\alpha}_{m}\cap\{g^{[\nu_{\alpha}(g)]}=...=g^{(k)}=0\}.$
With this in mind, we can sum up the purpose of this section in Table 2 which deals with how to determine the irreducible components of $\mathcal{C}^{0}_{m}$. Whenever $\mathcal{H}^{\gamma}_{m}$ is mentioned we impose the condition $\nu_{\gamma}(f)>m$ and whenever
@@
corresponds to a branch defined by a factor $f_{i}$ of $f$ with $\alpha\in T(f_{i})$, we impose the condition $m\geq\nu_{\alpha}(f)$ and $a=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$.
AA $B$ when does $A\subseteq B?$ reason behind $A\subseteq B$
@@
$\mathcal{H}^{\gamma}_{m}$ $\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,a}}$ $\gamma=\alpha+(1,1)$, $m=\nu_{\alpha}(f)$ this holds when $T(f)=T(f_{i})$
@@
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,a}}$ $\overline{\mathbf{A}^{2}_{m}\cap F^{\beta}_{j,b}}$ $\alpha=\beta$ and $f_{i}=f_{j}$ compare codimensions
@@
Lemma 4.
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve singularity, $\alpha\in T(f)$, with $\nu_{\alpha}(f)\leq m$. Then, there is no $(1,1)\leq\lambda<\alpha$ with $\nu_{\lambda}(f)>m-1$ i.e. $\mathcal{H}^{\lambda}_{m-1}\subseteq\mathcal{C}^{0}_{m-1}$.
Proof.
Suppose such $\lambda$ exists. Then, $\nu_{\lambda}(f)\geq m\geq\nu_{\alpha}(f)$. However, as $\lambda<\alpha$ then $\nu_{\lambda}(f)\leq\nu_{\alpha}(f)$ and $m=\nu_{\alpha}(f)=\nu_{\lambda}(f)$. Let’s set $H=\mathcal{H}^{\lambda+\varepsilon}_{m}$ and $V_{i}=\mathcal{H}^{\alpha}_{m}\cap\{f^{[\nu_{\alpha}(f_{i})]}_{i}=0\}$ where $f_{i}$ is an irreducible factor of $f$ with $\alpha\in T(f_{i})$, and
$\mathcal{H}^{\alpha}_{m-1}\subsetneq\mathcal{H}^{\lambda}_{m-1}\subseteq\mathcal{C}^{0}_{m-1},$
where $\varepsilon\in\{(1,0),(0,1)\}$ and $\pi_{m,m-1}:\mathcal{C}^{0}_{m}\rightarrow\mathcal{C}^{0}_{m-1}$ is the truncation morphism. By considering the preimage $\pi^{-1}_{m,m-1}$, we obtain
$V_{i}\subsetneq\mathcal{H}^{\alpha}_{m}\cap\{f^{(m)}=0\}\subseteq H.$
[…]
$(x^{(0)},...,x^{(\lambda_{1}-1)},y^{(0)},...,y^{(\lambda_{2}-1)})=\mathcal{I}_{H}=\{r\in\mathcal{O}_{\mathbf{A}^{2}_{m}}:r\cdot\mathcal{I}_{V_{i}}\subseteq\mathcal{I}_{H}\}$
is an equality for the defining ideals $\mathcal{I}_{H}=(\mathcal{I}_{H}:\mathcal{I}_{V_{i}})$ (colon ideal). By denoting $r=x^{(\lambda_{1}-1)}$ and $g=in_{\alpha}(f_{i})(x^{(\alpha_{1})},y^{(\alpha_{2})})\in\mathcal{I}_{V_{i}}$ in $r\cdot g\in\mathcal{I}_{H}$, then we can write $r\cdot g=\sum_{i=0}^{\alpha_{1}-1}x^{(i)}X_{i}+\sum_{j=0}^{\alpha_{2}-1}y^{(j)}Y_{j}$. Moreover, by annihilating the variables $x^{(0)},...,x^{(\lambda_{1}-2)},y^{(0)},...,y^{(\alpha_{2}-1)}$, we get
$x^{(\lambda_{1}-1)}[{y^{(\alpha_{2})}}^{p}-c_{i}\cdot{x^{(\alpha_{1})}}^{q}]=\sum_{i=\lambda_{1}-1}^{\alpha_{1}-1}x^{(i)}X^{\prime}_{i}.$
Now, we distinguish the following 2 cases.
  • If $\alpha_{1}>\lambda_{1}$ then by annihilating $x^{(\alpha_{1})},x^{(\alpha_{1}+1)},...,x^{(m)},y^{(\alpha_{2}+1)},...,y^{(m)}$, we notice that
    $x^{(\lambda_{1}-1)}{y^{(\alpha_{2})}}^{p}=\sum_{i=\lambda_{1}-1}^{\alpha_{1}-1}x^{(i)}X^{\prime\prime}_{i}.$
    So annihilating the variables $x^{(\lambda_{1})},...,x^{(\alpha_{1}-1)}$ leads to a contradiction because the non-zero monomial $x^{(\lambda_{1}-1)}{y^{(\alpha_{2})}}^{p}=0$ in a power series ring.
  • If $\alpha_{1}=\lambda_{1}$ then $\alpha_{2}>\lambda_{2}$. By taking $r^{\prime}=y^{(\lambda_{2}-1)}$ and $g^{\prime}=g$, we can write
    $r^{\prime}g=y^{(\lambda_{2}-1)}\cdot[{y^{(\alpha_{2})}}^{p}-c_{i}{x^{(\alpha_{1})}}^{q}]=\sum_{i=0}^{\alpha_{1}-1}x^{(i)}U_{i}+\sum_{j=0}^{\alpha_{2}-1}y^{(j)}V_{j},$
    which after annihilating the variables
    $x^{(0)},...,x^{(\alpha_{1}-1)},y^{(j)}:j\neq\lambda_{2}-1$
    leads to $c_{i}\cdot y^{(\lambda_{2}-1)}{x^{(\alpha_{1})}}^{q}=0$ (in a power series ring) which can not be the case.
Therefore, in both cases we arrived at a contradiction. Thus, the proof is complete. ∎
@@
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve singularity such that $T(f)=T(f_{i})$ where $f_{i}$ defines a branch of $f$. Let $\alpha,\beta\in T(f)\cap\mathbf{N}^{2}$ and $\gamma=\beta+(1,1)$ and suppose $\nu_{\gamma}(f)>m\geq\nu_{\alpha}(f)$. Then, we have
@@
if and only if $\alpha=\beta$ and $m=\nu_{\alpha}(f)$. Moreover, the inclusion on the left is always strict.
@@
Suppose first that $\alpha=\beta$ and that $\mathcal{H}^{\gamma}_{m}=\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$. Then,
$U:=F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}\subset\mathcal{H}^{\gamma}_{m}.$
However, $x^{(\alpha_{1})}\neq 0$ in $U$ and $x^{(\alpha_{1})}=0$ in $\mathcal{H}^{\gamma}_{m}$, which is a contradiction. So the inclusion if it occurs is always strict. Now, let us show the equivalence in the theorem. Suppose that $m=\nu_{\alpha}(f)$ and $\alpha=\beta$, then
@@
with $f^{[\nu_{\alpha}(f_{i})]}_{i}=[{y^{(\alpha_{2})}}]^{p}-c_{i}\cdot[{x^{(\alpha_{1})}}]^{q}\mod\alpha$ where $T(f_{i})=\mathbf{R}_{\geq 0}(p,q)$. So that we have the inclusion $\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$, with $m=\nu_{\alpha}(f)$ and $\gamma=\alpha+(1,1)$.
@@
To prove the other direction of the equivalence, suppose that we have the inclusion above with $m\geq\nu_{\alpha}(f)$ together with the condition $\alpha=\beta$. Then,
$\mathcal{H}^{\gamma}_{m}\neq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}.$
By comparing the codimensions of these irreducible sets:
$|\gamma|=|\alpha|+2>|\alpha|+1+m-\nu_{\alpha}(f),$
we arrive at the inequality $1>m-\nu_{\alpha}(f)\geq 0$ which shows that $m=\nu_{\alpha}(f)$.
@@
Now, assume on the other hand that $\alpha\neq\beta$ with $\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$. Let us recall that for every reduced plane curve $\mathcal{C}=\{g=0\}\subset\mathbf{A}^{2}$,
$m<\nu_{\gamma}(g)\Leftrightarrow(x^{(0)},y^{(0)},g^{(j)})_{0\leq j\leq m}\subseteq(x^{(i)},y^{(j)})_{0\leq i<\gamma_{1}\>\ 0\leq j<\gamma_{2}}.$
As $\mathcal{H}^{\gamma}_{m}\subseteq\mathcal{H}^{\alpha}_{m}\cap\{f^{(j)}_{i}=0\}_{0\leq j\leq a}\subseteq\mathcal{C}^{0}_{i,a}$, then
$(x^{(0)},y^{(0)},f^{(j)}_{i})_{0\leq j\leq a}\subseteq(x^{(i)},y^{(j)})_{0\leq i<\gamma_{1}\>\ 0\leq j<\gamma_{2}}$
and $\nu_{\gamma}(f_{i})>a=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$. If $a+h=\nu_{\gamma}(f_{i})-1$ then
$(x^{(0)},...,x^{(\alpha_{1}-1)},y^{(0)},...,y^{(\alpha_{2}-1)},f_{i}^{(j)})_{0\leq j\leq a+h}\subseteq(x^{(i)},y^{(j)})_{0\leq i<\gamma_{1}\>\ 0\leq j<\gamma_{2}},$
$\mathcal{H}^{\gamma}_{a+h}\subset\mathcal{H}^{\alpha}_{a+h}\cap\{f_{i}^{(j)}=0\}_{\nu_{\alpha}(f_{i})\leq j\leq a+h}$
is a strict inclusion. Therefore, $|\gamma|\geq 1+a+h-\nu_{\alpha}(f_{i})+1+|\alpha|$. By replacing $\gamma=\beta+(1,1)=b\cdot\alpha+(1,1)$, we find that
$b|\alpha|+2=|\beta|+2=|\gamma|\geq 2+|\alpha|+\nu_{\gamma}(f_{i})-1-\nu_{\alpha}(f_{i}),$
$b|\alpha|\geq|\alpha|+\nu_{\gamma}(f_{i})-1-\nu_{\alpha}(f_{i}).$
As $\nu_{\gamma}(f_{i})>a\geq\nu_{\alpha}(f_{i})$ then $b|\alpha|\geq|\alpha|$ and $b\geq 1$. We have $\mathcal{H}^{\beta}_{m-1}\subseteq\mathcal{H}^{\alpha}_{m-1}$ i.e. $\beta\geq\alpha$. Moreover, by the previous lemma, there is no $\lambda<\beta$ such that $\nu_{\lambda}(f)>m-1$. Thus, we must have $\alpha=\beta$.
Now suppose $m=\nu_{\alpha}(f)$. If
$\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})}\cap\mathbf{A}^{2}_{m}}=\mathcal{H}^{\alpha}_{m}\cap\{f_{i}^{[\nu_{\alpha}(f_{i})]}=0\}$
then again by comparing codimensions, (since the inclusion is strict), we get $|\gamma|>1+|\alpha|$. Writing $\beta=b\cdot\alpha$ (with $b>0$) in
$(b-1)|\alpha|=|\beta|-|\alpha|=|\gamma|-(2+|\alpha|)\geq 0,$
guarantees that $b\geq 1$ so that $\mathcal{H}^{\beta}_{m-1}\subseteq\mathcal{H}^{\alpha}_{m-1}$ i.e. $\beta\geq\alpha$. As there is no $\lambda<\beta$ such that $\nu_{\lambda}(f)>m-1$, we again must have $\alpha=\beta$. ∎
@@
Our next theorem deals with the fact that distinct infinite components (of the jet schemes of a Newton Plane Curves) come from distinct branches of this curve. This in turns means that whenever one such irreducible component is inside another infinite component, then those components are essentially the same.
@@
Before doing so, we need to recall the formula of the codimension of $G^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}$ (where $G$ is defined by an irreducible Newton non-degenerate $g$ that divides $f$) is given by:
@@
in which we set $k=\nu_{\alpha}(g)+m-\nu_{\alpha}(f)$. With this in mind, we have the following theorem.
@@
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve singularity. Let $f_{i}$ and $f_{j}$ define branches of $f$ such that
$T(f_{l})=\mathbf{R}_{\geq 0}\cdot\alpha_{l}\mbox{ with }\alpha_{l}\in\mathbf{N}^{2}$
and $\nu_{\alpha_{l}}(f)\leq m$ for $l\in\{i,j\}$. Set $k_{l}=\nu_{\alpha_{l}}(f_{l})+m-\nu_{\alpha_{l}}(f)$. If
@@
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{j}}_{j,k_{j}}}.$